The force of attraction between two masses m1 and m2, having dimensions that are small compared with their


separation distance R, is given by Newton’s third law, F=Gm1m2/R2, where G=6.67259x10-11 N-m2/kg2. What


is the total gravitational force that the sun (1.97x1030 kg) and the earth (5.95x1024 kg) exert on the moon (7.37x1022 kg) at


an instant when the earth, moon, and sun form a 90° angle? The earth-moon and sun-moon distances are 380x103 km


and) 150x106 km, respectively

1
Expert's answer
2021-11-06T01:35:18-0400

"F= \frac{Gm_1m_2}{R_2} = \frac{6.67259\times 10^{-11}\times 5.95\times 10^{24}\times 7.37\times 10^{22}}{380\times10^3}=77 \times 10^{30}"